OpenStudy (anonymous):

For denominators not equal to 0, simplify

OpenStudy (anonymous):



OpenStudy (anonymous):

@jdoe0001 @SolomonZelman

OpenStudy (anonymous):

normally try and factorise to eliminate factors so x^2 -1 = (x+1)(x-1)

OpenStudy (anonymous):

12x^2 -6x = 2x(x -3) interestingly enough, the one below is 3x-3 = 3(x-3)

OpenStudy (anonymous):

so at least on the right hand side you are left with 2x/3

OpenStudy (anonymous):

As far as I can see there is no simple factors to 2x^2 + x + 1 therefore answer is: (2x)(x^2 -1)/3(x^2+x+1) = 3x^3 -2x/3x^2+3x+3